OpenStudy (shelby1290):

A 6-m extension ladder leans against a vertical wall with its base 2m from the wall. How high up the wall does the top of the ladder reach? Round to the nearest tenth.

OpenStudy (shelby1290):



OpenStudy (juicstice):

just do Pythagorean theorem. x^2+2^2=6^2. can you solve for x?

OpenStudy (anonymous):

Yeah the Pythagorean theorem. a^2 b^2 = c^2 like what Juicstice said.

OpenStudy (anonymous):

*a^2 + b^2 = c^2 oops.

OpenStudy (shelby1290):

x^2+2^2=6^2 x^2+4=36 36-4=32 x^2=32 \[\sqrt{32}\] =5.6 x=6m

OpenStudy (juicstice):

not it would be 5.6

OpenStudy (shelby1290):

oh okay i thought i had to round it

OpenStudy (juicstice):

yea, to the nearest tenth

OpenStudy (shelby1290):

alright, so 5.6 stays the same. ok so then what's the next step

OpenStudy (juicstice):

you have it :).

OpenStudy (shelby1290):

oh really?! lol im stupid i thought i had to find the perimeter or something

OpenStudy (juicstice):

nope lolz

OpenStudy (shelby1290):

alrighty thank youuu